Evaluate limit 0to ln3 ∫ e^x(1+e^x)^1\2dx .


1
Expert's answer
2020-09-14T17:58:25-0400
"\displaystyle\int_{0}^{\ln3}e^x(1+e^x)^{{1 \over 2}}dx"

"\int e^x(1+e^x)^{{1 \over 2}}dx"

"u=1+e^x, du=e^xdx"


"\int e^x(1+e^x)^{{1 \over 2}}dx=\int u^{{1 \over 2}}du=\dfrac{2}{3}u^{{3 \over 2}}+C="

"=\dfrac{2}{3}(1+e^x)^{{3 \over 2}}+C="

"\displaystyle\int_{0}^{\ln3}e^x(1+e^x)^{{1 \over 2}}dx=\big[\dfrac{2}{3}(1+e^x)^{{3 \over 2}}\big]\begin{matrix}
 \ln3 \\
 0
\end{matrix}="

"=\dfrac{2}{3}(1+3)^{{3 \over 2}}-\dfrac{2}{3}(1+1)^{{3 \over 2}}="


"=\dfrac{16}{3}-\dfrac{4\sqrt{2}}{3}"